OpenStudy (anonymous):

gievn f(x)= 2x+1, h(x)= 2x^2 + 2x +9 and h(x) = f o g what is g(x)? please also explain what f o g is?

OpenStudy (anonymous):

\[f \circ g\]means f compsed of g. so let's look at your problem and apply it too get a better idea of exactly what is meant by that.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

\[f(x)=2x+1\text{ or }f(\,\,\, )=2(\,\,\, )+1\]so whatever goes in the parentheses on the left also goes in the parentheses on the right. when they want f composed of g, that means they want to know what f of g is, or what do you get when you put in the g function for x in f(x)... \[(f \circ g)(x) = f \left( g(x) \right)=2\left( g(x) \right)+1=2x^2+2x+9\] so we can kind of back out what g(x) is... \[(f \circ g)(x) =2\left( g(x) \right)+1=2x^2+2x+8+1=2(x^2+x+4)+1\] so \[g(x)= x^2+x+4\]

OpenStudy (anonymous):

i see so we basically break it down

OpenStudy (anonymous):

in this case, yeah